OpenStudy (anonymous):

I could use a little more help. I'm sure it is rather simple because I have worked some (with the help of others) that looked like this. Solve. 3x^2=33x-27

OpenStudy (anonymous):

\[3x ^{2}=33x-27\]

OpenStudy (anonymous):

try factoring out a 3, that should simplify the rest of it

jimthompson5910 (jim_thompson5910):

3x^2=33x-27 3x^2-33x+27 = 0 ... get everything to one side 3(x^2 - 11x + 9) = 0 ... factor out 3 x^2 - 11x + 9 = 0/3 ... Divide both sides by 3. x^2 - 11x + 9 = 0 Now use the quadratic formula to solve for x

OpenStudy (anonymous):

you should be able to solve for x now based on jim's working.

OpenStudy (whpalmer4):

You can also use the quadratic formula directly after Jim's second step: a = 3, b = -33, c = 27. Answers will be the same; arithmetic may be easier if you've factored out any common factors, however.